From the class survey, the mean birth weight of students is 7.1 pounds with a standard deviation of 1.1 pounds. At what weight is more than 80% of the students? Include a.) the distribution curve, b.) the z-curve, and c.) the solution.
Answer #1


P ( X > ? ) = 1 - P ( X < ? ) = 1 - 0.8 = 0.2
Looking for the probability 0.2 in standard normal table to calculate critical value Z = -0.84

-0.84 = ( X - 7.1 ) / 1.1
X = 6.176
P ( X > 6.176 ) = 0.8

Part b)